OpenStudy (anonymous):

olve the following system of equations. =+−3x8y−19 =+8x7y−6

OpenStudy (freckles):

hard to understand.

OpenStudy (freckles):

just so you know this is what you have typed: "olve the following system of equations. =+−3x8y−19 =+8x7y−6"

OpenStudy (anonymous):

-3x+8y=-19 8x+7y=-6